OpenStudy (anonymous):

Find all the real zeros of the function: f(x)=x^4-5x^3+7x^2+3x-10

OpenStudy (tkhunny):

Have you considered the factors of 10?

OpenStudy (anonymous):

Yes +-1,+-2,+-5,+-10 but I can't get from there I'm using synthetic division

OpenStudy (tkhunny):

Thaat's okay. That will find only the RATIONAL zeros. If they are not Rational, that is the right result.

OpenStudy (tkhunny):

On the other hand, you should try x = 2 and x = -1 again.

OpenStudy (anonymous):

But what I end up with is (x^3-6x^2) +(13x-10) and that doesn't factor out correctly

OpenStudy (tkhunny):

I don't know what that is. How did you come to that condition?

OpenStudy (anonymous):

I did synthetic division using -1

OpenStudy (tkhunny):

Please demonstrate. You should not end up with that.

OpenStudy (tkhunny):

 Well, there's the numbers for x = 2. I wish it had looked better.

OpenStudy (tkhunny):

I see. 1, -6, 13, -10 is good. Now do x = 2.

OpenStudy (anonymous):

Ok I tried 2 and I got that as we'll but then after you have to turn it into a poynomial equatiom

OpenStudy (anonymous):

*well

OpenStudy (tkhunny):

Did you get \(x^{2} - 4x + 5 = 0\)?

OpenStudy (anonymous):

No I got x^3 -3x^2 +x+5

OpenStudy (tkhunny):

We are going around in circles. x^4-5x^3+7x^2+3x-10 = (x-2)(x^3 - 3x^2 + x + 5) x^4-5x^3+7x^2+3x-10 = (x+1) (x^3 - 6x^2 + 13x-10) Pick one and do the other to it. x^4-5x^3+7x^2+3x-10 = (x+1)(x-2)(x^2 − 4x + 5)

OpenStudy (anonymous):

Ok thanks I guess that was easier